OpenStudy (anonymous):

Divide and simplify (x^4 y^4)/(root(3)(x^2 y))

OpenStudy (campbell_st):

ok... so you need index notation for the denominator \[(x^2y)^{\frac{1}{3}}\] which becomes \[x^{\frac{2}{3}}y^{\frac{1}{3}}\] so now you can use the index law for division of the same base, subtract the powers the problem now looks like \[\frac{x^4y^4}{x^{\frac{2}{3}}y^{\frac{1}{3}}}\] hope it helps

OpenStudy (anonymous):

Thank you! how would the answer look like ?

OpenStudy (campbell_st):

well it starts by \[x^{\frac{10}{3}}y^{\frac{11}{3}} = (x^{10}y^{11})^{\frac{1}{3}} = \sqrt[3]{x^{10}y^{11}}\]